17. 05. 2001

Language: ngerman

today  normal date format 17. Mai 2001
isodate  2001-05-17
numdate  17. 05. 2001
shortdate  17. 05. 01
TeXdate  2001/05/17
back to normal  17. Mai 2001
another day normal date format 30. Januar 1972
isodate  1972-01-30
numdate  30. 01. 1972
shortdate  30. 01. 72
TeXdate  1972/01/30
back to normal  30. Januar 1972

Language: naustrian

today  normal date format 17. Mai 2001
isodate  2001-05-17
numdate  17. 05. 2001
shortdate  17. 05. 01
TeXdate  2001/05/17
back to normal  17. Mai 2001
another day normal date format 30. Jänner 1972
isodate  1972-01-30
numdate  30. 01. 1972
shortdate  30. 01. 72
TeXdate  1972/01/30
back to normal  30. Jänner 1972

Language: english

today  normal date format 17th May 2001
isodate  2001-05-17
numdate  17/05/2001
shortdate  17/05/01
TeXdate  2001/05/17
back to normal  17th May 2001
another day normal date format 30th January 1972
isodate  1972-01-30
numdate  30/01/1972
shortdate  30/01/72
TeXdate  1972/01/30
back to normal  30th January 1972

Language: american

today  normal date format May 17, 2001
isodate  2001-05-17
numdate  05/17/2001
shortdate  05/17/01
TeXdate  2001/05/17
back to normal  May 17, 2001
another day normal date format January 30, 1972
isodate  1972-01-30
numdate  01/30/1972
shortdate  01/30/72
TeXdate  1972/01/30
back to normal  January 30, 1972

Changed ISO output

2000-12-24 2000–12–24 2000·12·24